account that started with $4,000 10 years ago and which now has grown to $7,000 You have an a. What annual rate of return have you earned (you have made no additional contributions to the account)? b.If the in account earns 17% per year from now on, what will the accounts value be 10 years from now? a. What annual rate of return have you earned (you have made no additional contributions to the account)? Your annual rate of return is[1%. (Round to two decimal places)
Answer #1

In the first question, interest rate is hidden or we can say we have to find the interest rate.

we have to guess the rate by hit and trial approach, in other words, we have to guess the rate or rates and see if it fits the equation or not.

I've calculated and come to the conclusion that if we take the interest rate 5.756% then we can solve the equation.

here is the equation.,

FV = PV*(1+i)^n

where,

FV = future value which is $7000

PV = present value which is $4000

i = interest rate, which we have to find by guess

and n = time period, which is 10 years

So, the equation will be,

$7000 = $4000*(1+i)^10

so if we put 5.756% as interest rate then we can equate this equation.

FV = $4000*(1+0.05756)^10

      = $7000.20 OR $7000

SO, the annual rate of return would be 5.756% or 5.76%.

b answer.

FV = PV*(1+i)^n

FV = 7000*(1+0.017)^10

FV = $33,647.80

So, 10years from now, the account value will be $33,647.80